OpenStudy (anonymous):

One cup of fresh cooked broccoli contains about 45.22 calories. About how many calories are contained in 5.6 cups of broccoli? Round your answer to the nearest hundredth. A.80.83 calories B.25.32 calories C.8.08 calories D.253.23 calories

OpenStudy (anonymous):

45.22*5=226.1 45.22*0.60=27.132 226.1+27.132=253.232